### v1

@@ Abstract
Automated market makers (AMMs) are a form of decentralized cryptocurrency exchanges and considered a prime example of Decentralized Finance (DeFi) applications. Their popularity and high trading activity have resulted in millions of on-chain transactions leading to serious scalability issues. In this paper, we address the on-chain storage overhead problem of AMMs by utilizing a new sidechain architecture as a layer 2 solution, building a system called $\mathsf{ammBoost}$. Our system reduces the amount of on-chain transactions, boosts throughput, and supports blockchain pruning. We devise several techniques to enable layer 2 processing for AMMs while preserving correctness and security of the underlying AMM. We also build a proof-of-concept of $\mathsf{ammBoost}$ for a Uniswap-inspired use case to empirically evaluate its performance. Our experiments show that $\mathsf{ammBoost}$ decreases the gas cost by 94.53% and the chain growth by at least 80%, and that it can support up to 500x of the daily traffic volume observed for Uniswap in practice.
@@ 1 Introduction
Cryptocurrencies and blockchain technology provide an innovative model that led to new applications and research frontiers, as well as reshaping the Internet and its digital services. Decentralized Finance (DeFi) is one of these applications in which blockchains are used to transform traditional financial services, that are usually centrally-managed, into fully decentralized ones. Many of these systems operate in open-access settings, thus removing market entrance barriers for customers, and enabling a transparent and intermediary-free interaction. Smart contracts strengthen this model by providing an automated way to negotiate contract terms and enforce agreements.
@@
Automated market makers (AMMs) are considered a prime example of DeFi services [64]. They build a platform for automated token trading by establishing liquidity pools for trading token pairs. An AMM is usually implemented as a dApp; a set of smart contracts on top of a smart contract-enabled blockchain, where Ethereum is the dominant choice so far. AMMs in general include operations for trading, called swaps, and other for managing liquidity and fee collection including mints, burns, and collects. Many popular AMMs are deployed in practice and widely used. Examples include Uniswap [26], Curve [6], DODO [7], and Sushiswap [21], which during the first half of 2023 commanded around 62-71%, 6.38-14.01%, 4.04-7.16%, and 0.69-3.67% of the top ten AMMs market share, respectively [1]. During the same period, the AMM industry has a total monthly trading volume of \$46 - \$95 billion, and an estimated total market cap of nearly \$16 billion as of January 2024 [23].
@@
Challenges. The popularity and high trading activity of AMMs led to efficiency problems. AMMs produce a massive number of on-chain transactions. On the one hand, this increases the underlying blockchain size or storage overhead, and on the second hand, it incurs large (gas) fees. This large workload does not only amplify storage cost, but also transaction processing/confirmation delays due to the low throughput of blockchains (Ethereum’s throughput is around 12 transaction/sec on average [9]).
@@
Concretely, and based on the traffic analysis that we conducted for 2023 (see Appendix 0.C), users of Uniswap V3 produced 20 million transactions on Ethereum in 2023 (the size of these transactions range from $400$ bytes to $3000$ bytes). This translates to adding around $20.2$ GB to the Ethereum blockchain. In its year of deployment (Nov 2018), Uniswap V1 generated 34,000 transactions, and in 2023, Uniswap’s various versions on Ethereum generated around 80 million transactions, leading to 231,7% increase in transaction volume. These numbers indicate that the scalability problem of AMMs is amplified over the years.
@@
Limitations of prior work. Improving blockchain scalability is an active research area. Many solutions have been proposed; some of them target layer 1 improving on the consensus itself, such as sharding [41, 66, 51], while others target layer 2 allowing for some form of off-chain processing while logging only state changes on the blockchain, such as payment channels and networks [42] and rollups [15, 30]. However, when it comes to AMMs, applying these solutions impact performance and security, and may not even reduce the storage cost. In sharding, localized workload division policies are used to reduce cross-shard transactions. For smart contract-enabled blockchains, this means that a dApp will be contained in one shard [32, 61], so the whole AMM will be on one shard. Thus, parallel processing among shards is not utilized. Others [57] attempt to shard the AMM by distributing liquidity pools among several shards, but still reliance on locked cross-shard transactions may degrade the AMM performance. Furthermore, in all sharding solutions all transactions are logged on-chain, so they do not cut the storage cost.
@@
Layer 2 solutions that allow computations, i.e., beyond just payment transfer as in payment networks [42], have limitations. Optimistic rollups have long contestation periods that may reach one week as in Optimism and Arbitrum [15, 48]. Moreover, they have security issues—verifiers could be centralized trusted entities [60], while incentive compatibility of non-trusted verifiers is still an open question [53, 4, 16] which may lead to adopting incorrect ledger state changes. Zero knowledge (ZK)-rollups [20, 38, 37] are costly; proof generation may take several minutes and it becomes worse when attesting to complex transactions. They also may have long transaction confirmation delay that may reach 24 hours as in zkSync Era [30]. Not to mention that many of these systems require a trusted setup.
@@
Sidechains [35, 45, 46, 49] are another instance of layer 2 solutions that can improve scalability. Despite their potential, existing efforts mostly focus on two-way peg, i.e., currency transfer between the sidechain and the mainchain, and all of them considered independent sidechains. That is, each chain has its own transactions, miners, and tokens. Such Independence and the focus on two-way peg limit the performance gains that can be achieved, and do not allow for workload sharing between the chains. Moreover, none of these solutions allow pruning stale records. [^1] Polygon [17] operates a fast (block time is around 2 sec) EVM-compatible sidechain. The whole AMM is run on this sidechain and tokens can be transferred to Ethereum, if desired, using bridges or atomic swap techniques. While polygon can benefit from a decrease in latency, by virtue of running a fast blockchain, it suffers from the same problems outlined above. Lastly, isolating an AMM on an independent sidechain impacts composability with other dApps running on Ethereum.
@@
A new approach. Sidechains seem to have potential in building an effective layer 2 solution to promote scalability. This has been observed in [54], who proposed a framework called chainBoost with a new sidechain architecture that permits workload sharing and arbitrary data exchange between the side and main chains, as well as blockchain pruning. chainBoost targets resource markers—Web 3.0 systems that offer decentralized digital services on top of the currency exchange medium, e.g., Filecoin [10] and Livepeer [14]. chainBoost directs all heavy/frequent service-related traffic to the sidechain, which in turn processes this traffic and produces concise summaries of the state changes that are used to sync the mainchain. Once these summaries are confirmed on the mainchain, the temporary blocks containing the actual traffic on the sidechain will be pruned. The empirical results in [54] show substantial performance gains in terms of blockchain size, transaction confirmation delays, and throughput. All of these are achieved without compromising security and while keeping the mainchain as the single truth of the system state.
These advantages motivated us to explore the following: Can we apply the chainBoost framework to AMMs and achieve similar performance gains and security guarantees?
@@
Contributions. In this paper, we answer this question in the affirmative and propose $\mathsf{ammBoost}$; a storage control and throughput boosting solution for AMMs. $\mathsf{ammBoost}$ introduces a novel approach for dividing the AMM functionality into two modules: one that resides on the mainchain (i.e., the underlying smart contract-enabled blockchain) and another that is operated by the sidechain. In particular, $\mathsf{ammBoost}$ offloads processing most transactions (including swaps, mints, collects, and burns) to the sidechain, and minimizes the functionality remaining on the mainchain. The latter is encapsulated in a base smart contract called $\mathsf{TokenBank}$, that manages the actual tokens that are being traded by tracking only the transaction summaries produced by the sidechain. It also includes all operations that must happen in real time, and so must be handled by the mainchain, such as flash loans.
@@
$\mathsf{ammBoost}$ solves several challenges related to applying sidechains to AMMs. The chainBoost framework assumes a mutual-dependency relation between the mainchain and the sidechain. That is, both chains operate in the same domain and have the same transaction format, services, and miner population. This is not the case for $\mathsf{ammBoost}$. The AMM is merely a dApp deployed at the application layer, so it does not modify how the mainchain protocol works. Also, the mainchain miners do not maintain the sidechain as in chainBoost, even they might not be aware of its existence. However, $\mathsf{ammBoost}$’s sidechain is impacted by the mainchain since the tokens and the AMM state (and some core functionalities) are on the mainchain. Thus, $\mathsf{ammBoost}$ introduces a unidirectional dependency relation: the sidechain is impacted by the mainchain but not vice versa. Furthermore, the sidechain is supposed to process all trading activities while the actual tokens reside on the mainchain.
@@
We resolve these challenges by introducing several techniques. First, we require the AMM to have its own miners to maintain the sidechain. So, like any blockchain, these miners are assumed to have a mining power with honest majority, they need to build a Sybil-resistant identity by, e.g., using a proof-of-stake approach, and they will be rewarded for maintaining the sidechain using, e.g., the AMM native token. Second, we introduce epoch-based deposits, where a user has to deposit on the mainchain the anticipated amount of tokens needed to back up her activities (or issued transactions) during an epoch on the sidechain. Third, we introduce pool snapshot-based and delayed token-payout trading. That is, the pool token balances are retrieved from the mainchain at the beginning of the epoch, which are used to compute swap prices processed on the sidechain. These balances evolve during the epoch based on transactions processed during that epoch. Users, however, do not get their actual tokens resulting from the processed transactions immediately. They have to wait until the end of the epoch once the $\mathsf{TokenBank}$ that resides on the mainchain is synced. At that point, the mainchain AMM state is updated, tokens are paid out to users, and any leftover in the epoch deposits is refunded to their owners.
@@
We analyze the security of $\mathsf{ammBoost}$ showing that it preserves the correct and secure operation of the underlying AMM. We also build a proof-of-concept implementation for a Uniswap-inspired use case, and conduct experiments to empirically evaluate the performance gains, in terms of blockchain size, confirmation delay, and throughput, that $\mathsf{ammBoost}$ can achieve. Our experiments show that $\mathsf{ammBoost}$ achieves a 94.53% gas cost reduction and 80.25% chain growth reduction (when compared to a Uniswap version deployed on Sepolia). Furthermore, they show that $\mathsf{ammBoost}$ can support large traffic volume, on the order of up to 500x of Uniswap’s daily transaction volume.
@@
Although the focus of this work is on controlling the storage overhead of AMMs and boosting their throughput, we believe that $\mathsf{ammBoost}$ may have far reaching impact. Its paradigm can enable more optimizations for AMMs, e.g., integration of privacy-preserving techniques via the sidechain. Also, $\mathsf{ammBoost}$ could be beneficial for other DeFi applications, and dApps in general, as it introduces a framework for operating application-specific sidechains interacting with smart contract-enabled blockchains.
@@ 2 Background
In this section, we provide an overview of the general functionality of AMMs and the chainBoost framework that we use in the design of $\mathsf{ammBoost}$.
Automated market makers. AMMs build platforms for token trading powered by the users themselves. This is done by establishing liquidity pools such that a pool trades a pair of tokens, say tokens $A$ and $B$, against each others. Users are divided into: clients which could be sellers and buyers, and liquidity providers (LPs). Providing liquidity comes from the sellers themselves since buying token $A$ requires paying the price using token $B$ (and vice versa), and from LPs who deposit tokens in the pool and obtain fees or rewards in return.
@@
At a basic level, an AMM implementation supports several transaction types. Among them there are exact output and input swaps, which allow users to specify the exact method to process their trades. For liquidity management, mints, burns and collects allow LPs to submit liquidity positions, collect their fees, and edit or even withdraw these positions, respectively. AMMs may provide additional services, such as flash loans [63] allowing clients to take advantage of arbitrage opportunities across different platforms. Furthermore, more sophisticated liquidity approaches are being adopted, e.g., concentrated liquidity [43]. It allows LPs to define the price range on which they desire their liquidity to be applied. The goal is to solve the inefficient use of provided funds in the older schemes that evenly distribute the provided liquidity across the pool’s price curve.
The functionality of an AMM is commonly implemented as a set of smart contracts on top of a smart contract-enabled blockchain, where Ethereum is the dominant choice so far. These contracts create and manage the liquidity pools, and provide the API needed to interact with the AMM. Residing on a public blockchain led to several financial and security issues [36, 55, 58], e.g., front-running attacks, sandwich attacks, miner/maximal extractable value, etc. Understanding and solving these issues are active research areas. We do not discuss these issues further since they are not the focus of this work; we target the on-chain storage cost and throughput of AMMs.
@@
As shown in Figure 1, the sidechain is managed by the mainchain miners, where for each epoch, a committee is elected to process the sidechain traffic during that epoch. The rest of the mainchain miners, who are not on the committee, do not process the sidechain traffic, thus reducing their load. To speedup agreement, chainBoost employs a practical Byzantine fault tolerance (PBFT)-based consensus (similar to those in [50, 47]) for the sidechain. [^2] Similar to chainBoost, to simplify the presentation, we adopt a leader-based PBFT in which a leader proposes a block for the committee to agree on (as in [50]. Nonetheless, voting-based PBFT (as in [47]) in which a set of proposers can propose blocks and voters will choose one of them can be used instead.
@@
As shown in Figure 1, the sidechain is composed of two types of blocks: temporary meta-blocks and permanent summary-blocks. For each sidechain round, the committee mines a meta-block containing the transactions they processed, so that once a transaction appears in a meta-block it is considered final. In the last round of the epoch, this committee mines a summary-block summarizing all state changes induced by the meta-blocks within that epoch. After that, it issues a sync-transaction containing the summarized state changes, which the mainchain miners use to update the relevant state variables on the mainchain. Once the sync-transaction is confirmed on the mainchain, all meta-blocks used to produce the respective summary-block are discarded. This significantly reduces the sidechain size, and subsequently, the mainchain size. At the same time, having permanent summary-blocks allows anyone can verify the source of the state changes recorded on the mainchain.
@@
Applicability of chainBoost to our setting. Our setting is different from the one in [54]: First, the mainchain and sidechain miners are different. In $\mathsf{ammBoost}$, the mainchain miners belong to a smart contract-enabled blockchain, and the AMM is simply an application deployed on that blockchain. Thus, the sidechain must have its own miner population from which the committee is elected, and a technique, such as proof-of-stake, to mitigate Sybil attacks. Second, in chainBoost the two chains are mutually-dependent, i.e., their security and valid operation depend on each other. In $\mathsf{ammBoost}$, the dependence is unidirectional; interruptions on the mainchain impact the sidechain since the base contract that keeps track of the AMM state resides on the mainchain, but not vice versa. Sidechain interruptions will indeed lead to invalid state of the base AMM contract, but this contract resides on the application layer and does not impact the underlying mainchain or other deployed dApps. Third, in $\mathsf{ammBoost}$, the actual token pools reside on the mainchain, while the trading and liquidity-related activities are handled by the sidechain. Thus, a mechanism is needed to handle token payouts and deposits to enable accepting and processing only valid transactions. We address these issues in our system design in Section 4.

### v3

@@ Abstract
Automated market makers (AMMs) are a form of decentralized cryptocurrency exchanges that have attracted huge interest lately. They are considered a prime example of Decentralized Finance (DeFi) applications, a large category under Web 3.0. Their popularity and high trading activity have resulted in millions of on-chain transactions leading to serious scalability issues in terms of throughput and on-chain state size. Existing scalability solutions, when employed in the context of AMMs, are either ineffective due to their large overhead, or suffer from security and centralization issues.
In this paper, we address these challenges by utilizing a new sidechain architecture as a layer 2 solution, building a system called $\mathsf{ammBoost}$. Our system reduces the amount of on-chain transactions, boosts throughput, and supports blockchain pruning. We devise several techniques to enable layer 2 processing while preserving the correct and secure operation of AMMs. These include a functionality-split and layer 2 traffic summarization paradigm, an epoch-based deposit mechanism, and pool snapshot-based and delayed token-payout trading. We also build a proof-of-concept of $\mathsf{ammBoost}$ for a Uniswap-inspired use case to empirically evaluate performance. Our experiments show that $\mathsf{ammBoost}$ decreases the gas cost by 96.05% and the chain growth by at least 93.42%, and that it can support up to 500x of the daily traffic volume observed for Uniswap in practice.
@@ 1 Introduction
Cryptocurrencies and blockchain technology provide an innovative model that led to new applications and research frontiers, as well as reshaping the Internet and its digital services (under what is called Web 3.0). Decentralized Finance (DeFi) is a large category under Web 3.0 in which blockchains are used to transform traditional financial services, which are usually centrally managed, into fully decentralized ones. Many of these systems operate in an open-access model, thus removing market entrance barriers for customers, and enabling a transparent and intermediary-free interaction. Smart contracts strengthen this model by providing an automated way to negotiate contract terms and enforce agreements.
@@
Automated market makers (AMMs) are considered a prime example of DeFi services [71]. They build a platform for automated token trading by establishing liquidity pools for token pairs. An AMM is implemented as a decentralized application (dApp); a set of smart contracts on top of a smart contract-enabled blockchain—where Ethereum is the dominant choice so far, that support operations for trading and liquidity management, such as swaps, mints, burns, and collects. AMMs are a huge industry with a total monthly trading volume of \$46 - \$95 billion (during the first half of 2023), and an estimated total market cap of nearly \$16 billion as of January 2024 [24]. Many popular AMMs are deployed in practice and widely used. Examples include Uniswap [27], Curve [8], DODO [9], and Sushiswap [22], which during the first half of 2023 commanded around 62-71%, 6.38-14.01%, 4.04-7.16%, and 0.69-3.67% of the top ten AMMs market share, respectively [1].
@@
Challenges. At the same time, AMMs are a huge scalability problem. The popularity and high trading activity of AMMs led to serious efficiency problems since they produce a massive number of on-chain transactions. On the one hand, this increases the underlying blockchain size or storage overhead, and on the second hand, it incurs large (gas) fees. This large workload does not only amplify state storage cost, but also transaction processing/confirmation delays due to the low throughput of blockchains (Ethereum’s throughput is around 12 transaction/sec on average [11]). [^1] A dApp on-chain storage includes the state of the smart contract account of this dApp, and the transaction history recorded on the blockchain that produced this state. A misconception about on-chain storage cost is counting only the contract latest state while ignoring the permanent on-chain storage of these transactions.
@@
Concretely, based on the traffic analysis that we conducted for 2023 (see Appendix 0.C), Uniswap V3 users produced 20 million transactions on Ethereum in 2023 (transaction sizes range from $400$ bytes to $3000$ bytes). This translates to adding around $20.2$ GB to the Ethereum blockchain. In its year of deployment (Nov 2018), Uniswap V1 generated 34,000 transactions, and in 2023, Uniswap’s various versions on Ethereum generated around 80 million transactions, leading to 231,7% increase in transaction volume. These numbers indicate that the scalability problem of AMMs is amplified over the years. This does not only impact the AMM itself, but also other dApps deployed on the underlying blockchain. Such contention drives users to put in high transaction fees so that miners would prioritize their transactions.
The challenge is to handle this scalability problem without impacting security (by, e.g., employing trusted third parties or weakening the security of consensus, which introduces new threat vectors).
@@
Limitations of prior work. Improving blockchain scalability is an active research area. Many solutions have been proposed; some of them target layer 1 improving on the consensus itself, such as sharding [42, 73, 53], while others target layer 2 allowing for some form of off-chain processing, such as payment channels and networks [43] and rollups [16, 31]. However, when it comes to AMMs, applying these solutions impacts performance and security, and may not even cut the storage cost. In sharding, localized workload division policies are used to reduce cross-shard transactions. For smart contract-enabled blockchains, this means that a dApp (so the whole AMM) will be contained in one shard [33, 66], thus parallel processing among shards is not utilized. Others [60] use static analysis to shard dApps by splitting them into commuting functional units that can be executed in any order, and assigning each unit to a shard. This cannot work for AMMs since their (per pool) operation is sequential. Distributing liquidity pools among shards has been proposed in [62], but the reliance on locked cross-shard transactions (to support multi-swaps and arbitraging) may degrade the AMM performance. All these sharding solutions log all transactions on-chain (i.e., the shards), so they do not cut the storage cost. [^2] Ledger pruning for sharding has been proposed in [53, 67] but in the UTXO model, i.e., as in Bitcoin; it is not for the account model (as in Ethereum) that represents smart contract-enabled blockchains on top of which AMMs are deployed.
@@
Layer 2 solutions that allow computations, i.e., beyond just currency transfer as in payment networks [43], have limitations. Optimistic rollups have long contestation periods that may reach one week as in Optimism and Arbitrum [16, 49]. Thus, a user cannot act based on the submitted state changes immediately, but has to wait until the end of the contestation period to ensure that the submitted results are valid. Moreover, they have security issues—verifiers (who validate the submitted changes during the contestation period) could be centralized trusted entities [65], while incentive compatibility of non-trusted verifiers is still an open question [56, 5, 17] which may lead to adopting incorrect ledger state changes. Zero-knowledge (ZK)-rollups [21, 39, 38] are costly; proof generation may take several minutes and it becomes worse when attesting to complex transactions. They also may have a long transaction confirmation delay that may reach 24 hours as in zkSync Era [31]. This impacts transaction processing and confirmation delays, forcing the users to wait longer for their transactions to be finalized. Not to mention that many of the used ZK systems require a trusted setup. [^3] Another instance of layer 2 solutions resorted to employing a centralized settlement party that matches trades and generate ZK proofs to prove settlement correctness [50]. Thus, this solution is a form of ZK-rollups that employs a single settlement party.
@@
Sidechains [36, 46, 47, 51] are another type of layer 2 solutions that can improve scalability. Despite their potential, existing efforts mostly focus on two-way peg, i.e., currency transfer between the sidechain and the mainchain, and all of them are considered independent sidechains. That is, each chain has its own transactions, miners, and tokens. Such independence and the focus on two-way peg limit the performance gains that can be achieved, and do not allow for workload sharing between the chains. Moreover, none of these solutions allow pruning stale records. [^4] In parallel to these academic efforts, many industrial initiatives have explored the utility of sidechains [61, 6, 70, 2, 55, 30]. Most of them focused on independent sidechains and two-way peg. Cosmos [6] allows some form of data exchange via publishing events and commitments on the destination chain. This worsens the storage problem due to data duplicates, and there is no support for blockchain pruning. Other instances in practice resorted to operating the whole AMM on an independent sidechain, i.e., as a separate system from the underlying smart contract-enabled blockchain. For example, Polygon [18] operates a fast EVM-compatible sidechain running along with Ethereum. The whole AMM is run on this sidechain and tokens can be transferred to Ethereum using bridges or atomic swap techniques. Isolating an AMM on an independent sidechain impacts composability with other dApps running on the mainchain (which is Ethereum in this case), limits interaction with the mainchain to merely currency transfers, and complicates system design and its security due to the involvement of bridges. Not to mention that this solution just moves the on-chain storage cost to the sidechain, which will have scalability issues on its own as the AMM user population grows.
@@
A new approach. Sidechains seem to have (so far under-utilized) potential in building an effective layer 2 solution to promote scalability. This has been observed in [57], who proposed a framework called chainBoost with a new sidechain architecture that has a mutual-dependence relation with the mainchain, thus permitting workload sharing and arbitrary data exchange, as well as blockchain pruning. chainBoost targets resource markers—Web 3.0 systems that offer decentralized digital services, e.g., Filecoin [12] and Livepeer [15]. chainBoost directs all heavy/frequent service-related traffic to the sidechain, which in turn processes this traffic and produces concise summaries of the state changes that are used to sync the mainchain. Once these summaries are confirmed on the mainchain, the temporary blocks containing the actual transactions on the sidechain are pruned. The empirical results in [57] show substantial performance gains in terms of blockchain size, transaction confirmation delays, and throughput. All of these are achieved without compromising security and while keeping the mainchain as the single truth of the system state.
These advantages motivated us to explore the following: Can we control the state growth of AMMs, and boost their throughput, in a secure and low-overhead way, and without isolating the AMM on a separate blockchain, using dependent-sidechains?
1.1 Our Contributions
We answer this question in the affirmative and propose $\mathsf{ammBoost}$; a secure storage control and throughput boosting solution for AMMs. In particular, we make the following contributions.
@@ 1.1 Our Contributions
System design. $\mathsf{ammBoost}$ introduces a novel approach for dividing the AMM functionality into two modules: one that resides on the mainchain (i.e., the underlying smart contract-enabled blockchain) and another that is operated by the sidechain. In particular, $\mathsf{ammBoost}$ offloads processing most transactions (swaps, mints, collects, and burns) to the sidechain, and minimizes the functionality remaining on the mainchain. The latter is encapsulated in a base smart contract called $\mathsf{TokenBank}$, that manages the actual tokens by tracking only the transaction summaries produced by the sidechain. It also includes all operations that must happen in real-time on the mainchain, such as flash loans.
@@
$\mathsf{ammBoost}$ solves several challenges related to applying dependent sidechains to AMMs. The chainBoost framework assumes a mutual-dependency relation between the main and side chains; both operate in the same domain and have the same transaction format, services, and miner population. This is not the case for $\mathsf{ammBoost}$; the AMM is merely a dApp deployed at the application layer, so it does not modify how the mainchain protocol works. Thus, the mainchain miners do not maintain the sidechain as in chainBoost, even they might not be aware of its existence. However, $\mathsf{ammBoost}$’s sidechain is impacted by the mainchain since the tokens and the AMM state (and some core functionalities) are on the mainchain. This means means that $\mathsf{ammBoost}$ introduces a unidirectional dependency relation: the sidechain is impacted by the mainchain but not vice versa. Furthermore, the sidechain should process all trading activities without a custody of the actual tokens, and must ensure that only transactions for which issuing users own tokens on the mainchain are accepted.
@@
We resolve these challenges by introducing several techniques. First, we require the AMM to have its own miners to maintain the sidechain. So, like any blockchain, these miners are assumed to have a mining power with an honest majority, they need to build a Sybil-resistant identity by, e.g., using a proof-of-stake approach, and they will be rewarded for maintaining the sidechain using, e.g., the AMM native token. Second, we introduce epoch-based deposits, where a user has to deposit on the mainchain the anticipated amount of tokens needed to back up her activities (or issued transactions) during an epoch on the sidechain. Third, we introduce pool snapshot-based and delayed token-payout trading. That is, the pool token balances are retrieved from the mainchain at the beginning of the epoch, which are used to compute trade prices processed on the sidechain. These balances evolve during the epoch based on the processed transactions processed. Users can use newly accrued tokens in trading since all balances are tracked, but they cannot withdraw the actual tokens directly. This is because the sidechain does not hold actual tokens; token payouts and deposit leftover refunds happen at the end of the epoch when $\mathsf{TokenBank}$ is synced.
Security analysis. We analyze the security of $\mathsf{ammBoost}$ showing that it preserves the correct and secure operation of the AMM.
@@
Implementation and evaluation. We also build a proof-of-concept implementation for a Uniswap-inspired use case, and conduct experiments to empirically evaluate the performance gains, in terms of blockchain size, confirmation delay, and throughput, that $\mathsf{ammBoost}$ can achieve. Our experiments show that $\mathsf{ammBoost}$ achieves a 96.05% gas cost reduction and 93.42% chain growth reduction (when compared to a Uniswap version deployed on the Sepolia testnet). Our experiments demonstrate that $\mathsf{ammBoost}$ can support large traffic volumes, on the order of up to 500x of Uniswap’s daily transaction volume. We also study impact of various configuration parameters on the performance gains of $\mathsf{ammBoost}$.
@@
Although the focus of this work is on controlling the storage overhead and boosting throughput of AMMs, we believe that $\mathsf{ammBoost}$’s paradigm can enable more optimizations for AMMs, e.g., integration of privacy-preserving techniques. Also, $\mathsf{ammBoost}$ could be beneficial for other DeFi applications, and dApps in general, as it introduces a framework for operating application-specific sidechains interacting with smart contract-enabled blockchains. We leave exploring such directions as part of our future work.
@@ 2 Background
We provide an overview of the general functionality of AMMs and the chainBoost framework that we use in the design of $\mathsf{ammBoost}$.
Automated market makers. AMMs build platforms for token trading powered by the users themselves. This is done by establishing liquidity pools such that a pool trades a pair of tokens, say tokens $A$ and $B$. Users are divided into: clients which could be sellers and buyers, and liquidity providers (LPs). Providing liquidity comes from the sellers themselves since buying token $A$ requires paying the price using token $B$ (and vice versa), and from LPs who deposit tokens in the pool and collect fees in return.
@@
At a basic level, an AMM implementation supports several transaction types: for trading, there are (exact input and exact output) swaps, and for liquidity management, there are mints, burns and collects that allow LPs to submit liquidity positions, collect their fees, and withdraw these positions, respectively. AMMs may provide additional services, such as flash loans [69] allowing clients to take advantage of arbitrage opportunities across different platforms. Furthermore, more sophisticated liquidity approaches are being adopted, e.g., concentrated liquidity [44] that enable defining a price range over which liquidity will be applied to address issues related to inefficient use of provided funds.
The functionality of an AMM is commonly implemented as a set of smart contracts on top of a smart contract-enabled blockchain, where Ethereum is the dominant choice so far. These contracts create and manage the liquidity pools, and provide the API needed to interact with the AMM. Residing on a public blockchain led to several financial and security issues [37, 58, 63], e.g., front-running attacks, sandwich attacks, miner/maximal extractable value, etc. Understanding and solving these issues are active research areas. We do not discuss these issues further since they are not the focus of this work; we target the storage cost and throughput of AMMs.
@@
As shown in Figure 1, the sidechain is managed by the mainchain miners, where for each epoch, a committee is elected to process the sidechain traffic during that epoch. The rest of the mainchain miners, who are not on the committee, do not process the sidechain traffic, thus reducing their load. To speedup agreement, chainBoost employs a practical Byzantine fault tolerance (PBFT)-based consensus (similar to those in [52, 48]) for the sidechain. [^5] Similar to chainBoost, to simplify the presentation, we adopt a leader-based PBFT in which a leader proposes a block for the committee to agree on (as in [52]). Nonetheless, voting-based PBFT (as in [48]) can be used instead.
@@
The sidechain is composed of two types of blocks (as shown in the figure): temporary meta-blocks and permanent summary-blocks. For each sidechain round, the committee mines a meta-block containing the transactions they processed, so that once a transaction appears in a meta-block it is considered final. In the last round of the epoch, this committee mines a summary-block summarizing all state changes induced by the meta-blocks within that epoch. After that, it issues a sync-transaction containing the summarized state changes, which the mainchain miners use to update the relevant state variables on the mainchain. Once the sync-transaction is confirmed on the mainchain, all meta-blocks used to produce the respective summary-block are discarded. This significantly reduces the sidechain size, and subsequently, the mainchain size. At the same time, having permanent summary-blocks allows anyone can verify the source of the state changes recorded on the mainchain.
@@
Applicability of dependent-sidechains to AMMs’ setting. Our setting is different from the one in [57]: First, the mainchain and sidechain miners run different protocols. In $\mathsf{ammBoost}$, the mainchain miners belong to a smart contract-enabled blockchain, and the AMM is simply an application deployed on that blockchain. Thus, the sidechain must have its own miner population, to run its protocol, and a technique, such as proof-of-stake, to mitigate Sybil attacks. [^6] Indeed, a miner can choose to operate on both the mainchain and the sidechain. Still this miner runs two protocols, one for each chain, rather than one protocol. Second, in chainBoost the two chains are mutually-dependent, i.e., their security and valid operation depend on each other. In $\mathsf{ammBoost}$, the dependence is unidirectional; interruptions on the mainchain impact the sidechain since the base contract that keeps track of the AMM state resides on the mainchain, but not vice versa. Sidechain interruptions will indeed lead to invalid state of the base AMM contract, but this contract resides on the application layer and does not impact the underlying mainchain or other deployed dApps. Third, in $\mathsf{ammBoost}$, the actual token reside on the mainchain, while the trading and liquidity-related activities are handled by the sidechain. Thus, a mechanism is needed to handle token payouts and deposits to enable accepting and processing only valid transactions. Devising techniques to address these issues resemble the core novelty of $\mathsf{ammBoost}$ system design.